OpenStudy (accidentalaichan):

Solve for x: 2 over 3 (x − 4) = 2x I don't get how to do this. Thanks for helping me ^^

OpenStudy (mathstudent55):

Is this the equation? \( \dfrac{2}{3(x - 4)} = 2x \) Or is this the equation? \( \dfrac{2}{3}(x - 4 ) = 2x \)

OpenStudy (accidentalaichan):

the second one

OpenStudy (anonymous):

\[\frac{ 2 }{3 }(x-4)=2x\] You want to follow the order of operations: Parentheses Exponents Multiply Divide Add Subtract Since you can't simplify the Parentheses anymore multiply the fraction to both values in the parentheses. Post the equation you have at that point.

OpenStudy (accidentalaichan):

x - 6 = 3x?

OpenStudy (accidentalaichan):

but the answer for that is -3

OpenStudy (anonymous):

not quite you multiplied something wrong: \[\frac{ 3}{ 2 } * \frac{ 2 }{3} (x-4) = 2x * \frac{ 3 }{ 2}\] try again looking at that